Problem 3: A 40-mm thick AISI 1020 steel plate is sandwiched between two 10-mm thick 2024-T3 aluminum plates and compresses w**Use Shigley's Mechanical Engineering Design, McGraw-Hill textbook as a reference
Answer #1

Poblen Fsrem Table (A-31) Jhe heght of Nut is (H)= Folution 8.4 mm 26.5Smm ISmm 1Omm 4omm 10mm L> +H(40t10+o+84 L we Rosundind) Ad 78.5ymm2 2. Ad 1) rem Table .1) Faoma-17 At58mm2 KAd As Este AdtA 207 Steet h7854x 53 x207 -54 (16) 58 (44 kb 247.582 MDe 5mm2 t10mm d= lOmm Stel tuppes half DT= B+ 2(d)tam30 . DT 26.55mm Estal= 207gfa From e(8-20) Ested)x d 0.577 T (D)+(or-d)